OpenStudy (anonymous):

which of the term of the sequence. (x^2-y^2),(x+y),(x+y)/(x-y)........is x+y/(x-y)^9

hartnn (hartnn):

first term, n=1 (x+y)/(x-y)^-1 note the denominator exponent = -1 2nd term , n=2 (x+y)/(x-y)^2 so, when n=2, den. exponent =0 similarly, n=3, denominator exponent = 1 ...... continuing this way, when denominator exponent is 9, can u tell what n will be ?

OpenStudy (anonymous):

x+y/(x-y)^9=(x^2-y^2) [1/(x-y]^n-1 =x+y/(x-y)^9=(x^2-y^2) 1/(x-y)^n-1 =(x+y)/(x^2-y^2(x-y)^9=1/(x-y)^n-1 =(x+y)/(x-y)(x+y)(x-y)9=1/(x-y)^n-1 =10=n-1 n=11?

hartnn (hartnn):

N=11 IS CORRECT, YES :)